OpenStudy (anonymous):

Little help please? ∫tan^6(2x)dx

OpenStudy (anonymous):

expand and usub?

OpenStudy (anonymous):

\[\int\limits \tan^6(2x) dx\] right

OpenStudy (anonymous):

So substitute u = 2x, du = 2dx \[\frac{ 1 }{ 2 } \int\limits \tan^6(u) du\] I think then you'll have to use a reduction formula

OpenStudy (anonymous):

\[\int\limits \tan^nx dx = ((\tan^n-1)x)/(n-1) - \int\limits \tan ^{n-2} xdx\]

OpenStudy (anonymous):

Yeah that's it

OpenStudy (anonymous):

You'll have to use that twice, n = 6, and n = 4

OpenStudy (anonymous):

And then you can write tan^u = sex^2u-1

OpenStudy (anonymous):

sec*

OpenStudy (anonymous):

tan^u being the second part being integrated?

OpenStudy (anonymous):

\[\begin{align*}\int\tan^62x~dx&=\int\tan^22x\tan^42x~dx\\\\ &=\int(\sec^22x-1)\tan^42x~dx\\\\ &=\int\sec^22x\tan^42x~dx-\int\tan^42x~dx\\\\ &=\int\sec^22x\tan^42x~dx-\int\tan^22x\tan^22x~dx\\\\ &=\int\sec^22x\tan^42x~dx-\int(\sec^22x-1)\tan^22x~dx\\\\ &=\int\color{blue}{\sec^22x}\color{red}{\tan}^4\color{red}{2x}~\color{blue}{dx}-\int\sec^22x\tan^22x~dx+\int\tan^22x~dx\\\\ &=\int\sec^22x\tan^42x~dx-\int\sec^22x\tan^22x~dx+\int\sec^22x~dx\\ &\quad \quad-\int dx \end{align*}\] Substitute \(\color{red}u=\color{red}{\tan2x}\), which would give \(\color{blue}{du}=\color{blue}{2\sec^22x~dx}\), then you have \[\frac{1}{2}\int u^4~du-\frac{1}{2}\int u^2~du+\int\sec^22x~dx-\int dx\]